### v3

# PersonalAI: A Systematic Comparison of Knowledge Graph Storage and Retrieval Approaches for Personalized LLM agents

Abstract
Personalizing language models that effectively incorporating user interaction history remains a central challenge in development of adaptive AI systems. While large language models (LLMs), combined with Retrieval-Augmented Generation (RAG), have improved factual accuracy, they often lack structured memory and fail to scale in complex, long-term interactions. To address this, we propose a flexible external memory framework based on knowledge graph, which construct and update memory model automatically by LLM itself. Building upon the AriGraph architecture, we introduce a novel hybrid graph design that supports both standard edges and two types of hyper-edges, enabling rich and dynamic semantic and temporal representations. Our framework also supports diverse retrieval mechanisms, including A*, water-circle traversal, beam search and hybrid methods, making it adaptable to different datasets and LLM capacities. We evaluate our system on three benchmarks: TriviaQA, HotpotQA, DiaASQ and demonstrate that different memory and retrieval configurations yield optimal performance depending on the task. Additionally, we extend the DiaASQ benchmark with temporal annotations and internally contradictory statements, showing that our system remains robust and effective in managing temporal dependencies and context-aware reasoning.
Index Terms:
GraphRAG, Graph Traversal Approaches, Knowledge Graphs Generation, MultiAgency, Question Answering,
[…]
=-21pt
I Introduction
TRecent advances in large language models (LLMs) have sparked growing interest in personalized AI systems capable in adapting to users based on their interaction history. Central to personalization is the challenge of encoding, storing and retrieving relevant information over long time horizons in a manner that supports efficient reasoning and response generation. While Retrieval-Augmented Generation (RAG) has become a widely used solution, enhancing factual recall by appending retrieved content to prompts, it remains limited by its unstructured nature and weak support for semantic relationships across stored memories.
In this work, we introduce a flexible graph-based memory framework designed to overcome these limitations by enabling structured, customizable representations of long-term memory and supporting advanced reasoning capabilities. Unlike traditional RAG pipelines that rely on dense vector similarity over raw text chunks, our system supports multiple memory formats: nodes, knowledge triples, thesis statements, episodic traces and dynamically organizes them into a knowledge graph. This structure allows the agent to represent, update and access semantic and temporal relationships with far greater control and interpretability.
Equally important, our framework supports a pluggable retrieval interface with multiple traversal mechanisms, including three variations of A* search, WaterCircle traversal, BeamSearch and hybrid strategies to adapt retrieval behavior to task demands and model capacity. We demonstrate that different memory and retrieval configurations yield optimal performance on different benchmarks and LLM scales, highlighting the versatility of our approach.
We build our system on top of the AriGraph architecture [anokhin2024arigraphlearningknowledgegraph], originally developed for LLM agents in interactive text environments such as TextWorld. AriGraph continuously maintains a structured knowledge base by extracting triples from observations, pruning outdated or redundant facts and integrating episodic and semantic vertices into a unified graph. Our framework extends this foundation by allowing customization of both save (memorization) and search (retrieval) modules, supporting task-specific tuning and component evaluation.
In summary, our main contributions are as follows:
  1. 1.
    We present a highly flexible external memory architecture that can be tuned via orthogonal hyperparameters for storage and retrieval.
  2. 2.
    We propose and evaluate six retrieval over structured knowledge graph, achieving superior performance across various datasets compared to GraphRAG baselines.
  3. 3.
    We enhance the DiaASQ benchmark by incorporating temporal structures into dialogue representations and demonstrate that our framework can leverage such structures to improve temporal reasoning.
This work provides a general and extensible framework for integrating long-term memory and adaptive reasoning into LLM agents, advancing the state of personalized and context-aware language generation.
II Related Work
In recent years, substantial advancements have been made in open-domain question answering (QA) and the personalization of language models. Techniques leveraging Wikipedia as a broad knowledge source [chen-etal-2017-reading] have successfully incorporated large-scale machine reading, integrating document retrieval with textual comprehension. Furthermore, dense representation-based methods for passage retrieval [karpukhin-etal-2020-dense] have demonstrated superior efficacy compared to traditional sparse retrieval approaches, such as TF-IDF and BM25, particularly in scenarios with sufficient training data.
The development of dense retrievers has demonstrated significant progress, particularly through integration of contrastive learning with unsupervised settings, which has shown promising results across diverse scenarios and outperformed conventional approaches such as BM25 [robertson2009probabilistic] and Contriever [izacard2021contriever]. Concurrently, pre-trained language models incorporating non-parametric memory access have been proposed for knowledge-intensive tasks. A notable example is retrieval-augmented generation (RAG) models, which integrate parametric and non-parametric memory mechanisms to improve question-answering performance [lewis2021retrievalaugmentedgenerationknowledgeintensivenlp].
Recent advancements in unsupervised dense retrieval models, such as ART, have demonstrated the ability to achieve state-of-the-art performance while eliminating reliance on labeled training data [sachan-etal-2023-questions]. In the domain of knowledge graph-based approaches, frameworks such as GraphReader [li2024graphreaderbuildinggraphbasedagent]) incorporate structured reasoning mechanisms to facilitate knowledge extraction and representation, with a particular emphasis on enhancing long-context reasoning capabilities.
For personalized models, AriGraph [anokhin2024arigraphlearningknowledgegraph] introduces a framework that integrates episodic memory and long-term planning using knowledge graphs. Similarly, HippoRAG [gutiérrez2024hipporagneurobiologicallyinspiredlongterm] employs personalized algorithms to improve question-answering (QA) performance by constructing semantic graphs, demonstrating notable advancements over conventional extraction methods. MemWalker [chen2023walkingmemorymazecontext] and PEARL [sarthi2024raptorrecursiveabstractiveprocessing] address challenges associated with context length, proposing architectures capable to efficiently traversing and consolidating information from large-scale documents.
Additionally, ReadAgent [lee2024humaninspiredreadingagentgist] addresses the challenge of processing long-text contexts by structuring content into memory episodes. Meanwhile, KGP [wang2023knowledgegraphpromptingmultidocument] proposes Knowledge Graph Prompting, a method that enhances multi-document question answering by constructing knowledge graphs to improve contextual reasoning.
These advancements demonstrate a sustained emphasis on improving language models comprehension, retrieval, and personalization capabilities. Such progress facilitates the development of more sophisticated systems, which leverages knowledge graphs to deliver personalized and contextually enriched interactions.
III Methods
III-A Memory Structure
In this study, we employ a graph knowledge base as external memory to enhance a large language model’s (LLM) question-answering capabilities. Memory model $G=(V_{o},E_{o},V_{t},E_{t},V_{e},E_{e})$ consists of semantic $(V_{o},E_{o},V_{t},E_{t})$ and episodic $(V_{e},E_{e})$ memory vertices and edges. In turn, semantic vertices and edges are classified into theses and objects. To capture and structure information from weakly structured sources, specifically, unstructured natural language texts $d_{i}$, the construction of that memory is performing by LLM in automatic manner. This memory graph provides a comprehensive representation of the original texts and comprises the following elements (see Figure 1):
  1. 1.
    $V_{o}$ is a set of object vertices. Each object vertex represents atomic concept extracted from corresponding $d_{i}$.
  2. 2.
    $E_{o}$ is a set of object edges. Object edge is a tuple $(v,rel,u)$, where $v$, $u$ are object vertices and relrel is a text-attributed relationship between them, that captures its direct associations. Object edges essentially represent triples integrated in the semantic memory.
  3. 3.
    $V_{t}$ is a set of thesis vertices. Thesis vertex encapsulate complete atomic thought, expressed in the corresponding $d_{i}$.
  4. 4.
    $E_{t}$ is a set of thesis edges. Thesis edges serves as hyper-edges, linking corresponding set of object vertices, extracted from the same $d_{i}$ and belonging to the given $v^{j}_{t}$.
  5. 5.
    $V_{e}$ is a set of episodic vertices. Each episodic vertex corresponds to an original text passage ($v^{i}_{e}=d_{i}$) and serve as hyper-edge, linking corresponding/relative vertices.
  6. 6.
    $E_{e}$ is a set of episodic edges. Each episodic edge $e^{i}_{e}=(v^{i}_{e},V^{i}_{s})$ connects all semantic vertices $V^{i}_{s}$, extracted from $d_{i}$, with each other by the corresponding episodic vertex $v^{i}_{e}$.
Figure 1: Example of a graph fragment, constructed from natural language text using our method, with object (green), thesis (yellow) and episodic (blue) vertices
III-B Memory Construction
By the terminology defined above, the process of constructing a memory graph from weakly structured sources can be decomposed into three key steps: (1) formulating vertices and their edges, (2) generating hyper-edges using an LLM, and (3) parsing the LLM’s output to store extracted information in a structured format (e.g., as subject-relation-object triples). An LLM prompts for extracting semantic memories from textual sources is provided in Appendix A. Full memory construction pipeline (Memorize pipeline) is demonstrated on Figure 2.
Figure 2: High level architecture of proposed Memorize pipeline for LLM-based triples extraction from unstructured texts on natural language and memory construction
The identification and retrieval of outdated information within the memory are performed through the following procedure. First, the vertices present in the newly extracted triples are compared against existing vertices in the memory graph to detect matches. Upon identifying matching vertices, these serve as the initial set for a breadth-first search (BFS), which traverses all associated standart- and hyper-edges. Subsequently, a specialized prompt is employed to instruct LLM to update retrieved knowledge with newly extracted data (see Appendix B for corresponding LLM prompts). If any triple are successfully updated, the corresponding outdated instance are removed from memory model.
III-C Information Search in Memory
Information search pipeline (QA pipeline) for memory model is shown in Figure 3.
Figure 3: High level architecture of proposed QA pipeline for generating answers to the questions based on constructed memory graph
As illustrated in Figure 3, pipeline comprises four primary stages, with the third stage further subdivided into two substages. QA pipeline begins by accepting a natural language question as input, which is then processed by the Entities Extractor module to extract key entities from it. In the second stage, these entities are passed to the Entities to Vertices Matcher module, where they are aligned with corresponding/relevant vertices in the memory graph.
The third stage involves two sequential operations: retrieval and filtering of graph triples based on semantic similarity to the input question. First, the Memory Graph Triples Retriever module initiates a graph traversal algorithm, using the matched entities from the second stage as starting vertices to retrieve a set of candidate triples. Subsequently, the Triples Filter module ranks these triples by computing their semantic similarity to the question via vector embeddings, retaining only the top $N$ (a predefined hyperparameter) most relevant items.
Finally, in the fourth stage, the Conditional Answer Generator module synthesizes a natural language answer, conditioned on a retrieved/filtered triples. The output of QA pipeline is the generated answer in a string format. Used LLM prompts for described modules is provided in Appendix C.
This pipeline architecture is designed based on three key considerations. First, to achieve an accurate initial approximation of the relevant subgraph, it is essential to align the key entities in the input question with semantically similar vertices in the existing memory graph. The reasoning behind this is that the information necessary for generating a correct response is typically localized within the subgraph containing these key entities. Second, the triples extracted from the knowledge graph exhibit only weak conditioning on the input question, limiting their direct applicability. Third, large language models (LLMs) are constrained by a fixed maximum sequence length for processing inputs in a single inference step, necessitating efficient retrieval and subgraph selection strategies.
III-D Retrieval Algorithms
The primary role of the constructed memory graph in our proposed system is to enable the accurate retrieval of information required for responding to specific user questions. This process necessitates a careful optimization between two key criteria: relevance and completeness. Relevance ensures that all retrieved information is directly pertinent to the question, whereas completeness guarantees the inclusion of all necessary contextual data, even if some retrieved elements may be extraneous. Consequently, striking an optimal balance between these factors is essential for achieving efficient and effective knowledge extraction.
To accomplish this, we design and implement advanced retrieval algorithms capable of dynamically balancing the trade-off between completeness and relevance through configurable parameter settings. These algorithms constitute a critical component of the question-answering pipeline, systematically traversing the memory graph to aggregate information essential for generating precise and contextually appropriate responses.
A*. The A* algorithm is a widely used method for graph traversal, particularly valued for its ability to efficiently identify shortest paths between vertices in a graph. In the context of our question-answering pipeline, this algorithm extracts and retains triples encountered along these shortest paths while eliminating duplicates based on a content. We hypothesize that triples, obtained through this process, contains necessary information for generating accurate responses. For traversal, we treat our memory graph as unweighted and undirected, using a constant distance metric between adjacent vertices. To optimize pathfinding efficiency, we evaluate three distinct heuristics for the $h$-metric:
  1. 1.
    Inner Product (IP): This heuristic computes dot product of embeddings between current and target vertices.
  2. 2.
    Weighted Shortest Path: This approach scales inner product metric by the length of shortest path, determined via Breadth-First Search (BFS).
  3. 3.
    Averaged Weighted Shortest Path: This heuristic calculates average of inner product distances between adjacent vertices along the path from start vertex to current vertex, as well as the direct path from current vertex to target vertex, and further weights this value by BFS-derived shortest path length.
WaterCircles. This graph-based extraction method employs a breadth-first search (BFS) algorithm to retrieve relevant knowledge. Query entities are first mapped to their corresponding/relevant vertices in the memory graph. Algorithm begins with exploring adjacent vertices of initial vertices and iteratively expands to neighboring vertices in subsequent steps, constructing radiate pathways. When pathways originating from different starting vertices intersect, triples, formed at these intersections, are aggregated into a primary list, while all traversed triples are compiled into a secondary list. Algorithm ultimately returns a subset of triples, selecting $N$ from the primary list and $K$ from the secondary list, where $N$ and $K$ are configurable hyperparameters.
When the memory graph incorporates not only direct object relations (object triples), but also associations between objects and text fragments (referred to thesis and episodic triples), a modified breadth-first search (BFS) algorithm is employed as follows:
  1. 1.
    Traversal Initialization: The search begins at vertices, that matches entities from the input question.
  2. 2.
    Text Fragment Analysis: During traversal, identified text fragments are examined for occurrences of other question entities, distinct from the entity that originated the path to the given fragment. For each fragment, the number of detected entities, denoted as $N_{intersections}$, is computed.
  3. 3.
    Ranking Triplets: The list of thesis and episodic triplets is then sorted in descending order based on $N_{intersections}$.
This triplet extraction strategy aims to enhance relevance and accuracy in retrieving information from the memory graph, thereby bolstering the effectiveness of AI-driven question-answering systems.
BeamSearch. Given a starting vertex, algorithm constructs $N$ (a hyperparameter) semantically relevant paths in response to the input question. This approach is inspired by beam search, a token generation strategy commonly employed in large language model (LLM) inference. The resulting paths are consolidated into a single list, with duplicate triples removal based on their string content. Traversal process is governed by the following hyperparameters:
  1. 1.
    Max depth: The maximum allowable depth for path construction.
  2. 2.
    Max paths: The maximum number of paths to generate.
  3. 3.
    Same Path Intersection by Vertex: If enabled, a path may revisit a vertex; otherwise, vertex revisitation is prohibited.
  4. 4.
    Diff Paths Intersection by vertex: If enabled, distinct paths may share vertices; otherwise, vertex sharing is disallowed.
  5. 5.
    Diff Paths Intersection by Edge: If enabled, distinct paths may share edges; otherwise, edge sharing is forbidden.
  6. 6.
    Final sorting mode: Determines the method for filtering the final set of paths. The search yields two path subsets: (1) ended_paths – path terminated before reaching the depth limit and (2) continuous_paths – paths that reached the depth limit. Each path is assigned a relevance score. The selection process varies based on the chosen mode. If ended_first–value is specified, then: ended–paths will be sorted in descending order by relevance and the first max_paths–paths will be selected. If there are fewer ended–paths than max_paths–value, then continuous–paths will be sorted by relevance and the first $N$ missing paths will be selected from them. If continuous_first–value is specified, then paths will be selected in the same way as ended_first–value, only first sorting/selection by continuous–paths, and then by ended–paths. If mixed–value is specified, then ended– and continuous–paths will be combined into one list, sorted in descending order by relevance, and the first $N$ max_paths–paths will be selected from the resulting list.
Mixed Algorithm. This algorithm integrates A*, WaterCircles and BeamSearch strategies to enhance extraction efficacy. By combining these approaches, we ensures that triples not captured by one method (e.g., WaterCircles) may still be retrieved by another (e.g., BeamSearch), thereby improving overall recall. The final set of triples, which is subsequently passed to the LLM for answer generation, constitutes the union of outputs derived from A*, WaterCircles and BeamSearch algorithms.
By evaluating these diverse algorithms, this study highlights advancements in extracting pertinent information from knowledge graphs, thereby supporting the robust architectural framework required for personalizing responses in large language model (LLM) agents.
IV Experiment Set-Up
IV-A Datasets
To evaluate proposed retrieval algorithms, we conducted experiments across three distinct benchmarks: DiaASQ [li2023diaasqbenchmarkconversational], HotpotQA and TriviaQA. This selection was designed to evaluate our framework across varying domains, structural complexities and reasoning requirements.
The primary evaluation dataset is DiaASQ, which consists of user dialogues from a Chinese forum focused on mobile device characteristics. A key feature of this dataset is the inclusion of structured ”true statements” that encapsulate the core semantic content of each dialogue. We procedurally generated evaluation questions from these statements to ensure precise assessment. To further evaluate framework’s capability in handling temporal dynamics and contradictory information, critical for personalized agents, we extended DiaASQ with explicit temporal annotations and internally contradictory statements.
To ensure broad applicability and mitigate potential dataset bias and limited domain diversity, we supplemented DiaASQ with two widely-used, general-domain QA benchmarks:
  1. 1.
    HotpotQA: Selected for its requirement of multi-hop reasoning across multiple documents.
  2. 2.
    TriviaQA: Chosen for its factoid-style, open-domain questions that test broad knowledge retrieval.
These datasets provide complementary challenges, moving beyond a single domain (mobile devices) to perform evaluation on general world knowledge and complex reasoning.
Considering the computational and engineering complexity of constructing and traversing large memory graphs, we created manageable yet representative subsets from the original datasets. This step was necessary to enable the iterative experimentation required for tuning multiple retrieval algorithms and LLM configurations within practical resource constraints. The preprocessing involved filtering contexts by length and, for TriviaQA, segmenting long documents into coherent chunks. The resulting subsets used for knowledge graph construction and QA pipeline evaluation are summarized in Table I. Detailed preprocessing steps and dataset statistics are provided in Appendix D.
Dataset #qa-pairs #contexts
DiaASQ 4800 3483
HotpotQA 2000 3933
TriviaQA 500 4925
TABLE I: Characteristics of prepared datasets for QA pipeline evaluation
Using these benchmarks, memorization and question-answering functionality of our PersonalAI framework is systematically compared against existing Retrieval-Augmented Generation (RAG) and GraphRAG baselines, as detailed in subsequent sections.
IV-B Models
For memory graph construction (Memorize pipeline) and information retrieval (QA pipeline) we evaluate a series of 7B/8B and large-scale ($\gg$ 14B) language models to assess their performance in handling these tasks. The selected models include Qwen2.5 7B, DeepSeek R1 7B, Llama3.1 8B, GPT4o-min, and DeepSeek V3. To generate vector representations (embeddings) of natural language text data, we employ a multilingual E5-large model [^1] https://huggingface.co/intfloat/multilingual-e5-small.
IV-C Graph-traversal Algorithms
Graph-traversal algorithms evaluation is conducted for A*, WaterCircles (WC) and BeamSearch (BS) as well as their combinations: ”WC + BS”, ”A* + BS”, ”A* + WC”. The values of hyperparameters for base algorithms were fixed (see Appendix E).
IV-D Graph-traversal restrictions
Additionally, we systematically varied the values of hyperparameters governing the traversal constraints applied to the graph during algorithm execution. These constraints determine which vertex types are excluded from traversal: ”E” prohibits traversal of episodic vertices, ”T” prohibits traversal of thesis vertices, and ”O” prohibits traversal of object vertices. The ”all” key word means that restrictions about graph vertices traversal is not applied.
IV-E Summary of experiment configurations
Each QA configuration was evaluated on 100 question-answer pairs from DiaASQ, HotpotQA, and TriviaQA datasets. The same LLM was used both for generating responses within a given QA configuration and for constructing the corresponding memory graph to execute the QA pipeline. Consequently, for each fixed dataset/model pairing, 22 distinct QA configurations were derived. In total, 308 QA configurations were evaluated (the QA pipeline was not executed on TriviaQA/GPT4o-mini configurations due to resource constraints).
IV-F Memory construction setting
Our implementation of memory consists of two main parts: graph part and vector part. The graph part stores textual representations of object, thesis and episodic vertices, their properties and relationships (edges) between them. Neo4j is used as the graph database for this part of the system. The vector part of memory stores vector representations (embeddings) of elements from the graph part to measure semantic similarity of texts for QA pipeline execution. Milvus is used as a vector database for this part of the system. Our memory model also implements a caching mechanism for storing intermediate results of QA pipeline components to reduce overall time spent to process incoming questions. This component utilizes two non-relational databases: Redis and MongoDB. During our experiments cache was enabled. All used databases were hosted and run on a single machine in personal docker containers. Medium size LLM (7B/8B) were used from local hosted Ollama docker container. LLM inference during memory construction and QA pipeline processing performed on a single NVIDIA TITAN RTX 24GB.
To evaluate QA pipeline we construct 14 memry graphs based on described datasets and LLMs above. Important to note, that we disable stage there we searching and deleting outdated knowledge in memory because of the experiments scope. An average speed of adding text fragments per minute to the memory for a given database configuration is 1̃.35 (with an average length of a processing texts 550-650 characters). Detailed characteristics of constructed memory graphs can be observed in Appendix G.
V Evaluation
Traditional statistical evaluation metrics such as BLEU [papineni2002bleu], ROUGE [lin2004rouge], and Meteor Universal [denkowski2014meteor] struggle to distinguish syntactically similar but semantically distinct texts. While semantic methods like BERTScore [zhang2019bertscore] were introduced to address these limitations, our experiments reveal that BERTScore lacks sufficient differentiability, often failing to capture nuanced distinctions between correct and incorrect answers. Therefore, we adopt the LLM as a judge [zheng2023judging] framework and choose Qwen2.5 7B. The judge evaluates QA pairs using a structured prompt containing the question, ground truth and model answer. It labels $1$ for correct answers and 0 for incorrect ones, and we use accuracy as our main metric. Corresponding LLM-prompts and details are provided in Appendix F.
Also, for comparison of proposed QA pipeline with existing RAG and GrapRAG methods the ExactMatch metric is calculated with ”ignore_case” and ”ignore_punctuation” hyperparameters set to True.
VI Experiments and Results
Based on experimental results, a comparative table, summarizing best-performing QA configurations by LLM-as-a-Judge metric, was compiled (see Table II).
LLM /
Dataset
Qwen2.5 7B DeepSeek R1 7B Llama3.1 8B GPT4o-mini DeepSeek V3
DiaASQ
0.22 / BS / all
0.12 / AS / E
0.19 / BS / E
0.5 / BS + WC / E
0.47 / BS + WC / O
HotpotQA
0.24 / BS / O
0.19 / BS / O
0.47 / BS / O
0.77 / BS + WC / all
0.76 / BS + WC / T
TriviaQA
0.34 / BS / E
0.27 / AS / E
0.66 / BS + AS / E
0.87 / BS + WC / all
Mean: 0.27 0.19 0.44 0.77 0.70
TABLE II: Best QA configurations ranked by LLM-as-a-Judge metric across all experiments. Corresponding cells contains LLM-as-a-Judge score, used retrieval algorithm and the type of restriction applied to the graph during traversal. Shortcuts for retrieval algorithms: BA – BeamSearch; AS – A*; BS + AS – hybrid of BeamSearch and A*; BS + WC – hybrid of BeamSearch and WaterCircles. Shortcuts for graph restrictions: all – no restrictions applied; E – episodic vertices excluded from traversal; T – thesis vertices excluded; O – object vertices excluded.
As shown in Table II, Qwen2.5 achieved best performance (0.27) among 7B models. Among all evaluated configurations, highest overall effectiveness (0.77) was reached by setups incorporating GPT-4o-mini. Notably, top-performing 7B configurations predominantly relied on BeamSearch, especially under constraint that traversal restricted through episodic vertices. In contrast, best DeepSeek V3 configurations frequently adopted a hybrid strategy combining BeamSearch and WaterCircles. Across high-performing configurations more broadly, BeamSearch consistently appeared as a key component within the retrieval pipeline.
To evaluate the effect of imposed constraints on the quality of the QA pipeline, we construct two distinct distributions of values, each corresponding to a specific model/dataset pair and retrieval algorithm. These distributions then were averaged across datasets and models for configurations exhibiting lowest and highest LLM-as-a-Judge values, as detailed in Table III.
Vertex types restrictions 7B 8B 14B+
worse configs best configs worse configs best configs worse configs best configs
E 3% 44% 9% 45% 27% 7%
T 84% 12% 64% 31% 20% 73%
O 13% 44% 27% 25% 53% 20%
TABLE III: Impact of various constraints, imposed during memory graph traversal, on the quality of QA pipeline: ”worse configs” – distribution for low quality configurations; ”best configs” – distribution for high quality configurations
As demonstrated in Table III, for 7B/8B models, majority ($\approx$74%) of configurations yielding the lowest response quality impose restrictions on traversing thesis-type vertices. Conversely, a significant proportion of high-quality configurations restrict traversal of episodic and object vertices ($\approx$44% and $\approx$34% respectively). This suggests that thesis-type memories contain critical information for generating relevant responses, whereas the inclusion of episodic and object memories introduces noise into the context, thereby degrading output quality. For larger-scale models, the trend differs: 53% of low-quality configurations restrict traversal of object vertices, while 73% of high-quality configurations restrict traversal of thesis-type vertices. This implies that larger models exhibit greater robustness in handling conditional generation from lengthy or noisy episodic memories, rendering thesis-based information redundant.
Retrieval algorithm 7B 8B 14B+ worse configs (w restr) best configs (w restr) other configs (w/o restr) worse configs (w restr) best configs (w restr) other configs (w/o restr) worse configs (w restr) best configs (w restr) other configs (w/o restr) WC 0.09 0.34 0.55 AS 0.1 0.175 0.14 0.29 0,36 0.41 0.23 0.36 0.33 BS 0.025 0.18 0.06 0.26 0.5 0.36 0.48 0.6 0.65 WC+BS 0.033 0.095 0.02 0.30 0.39 0.32 0.62 0.7 0.68 BS+AS 0.02 0.175 0.01 0.25 0.48 0.36 0.48 0.64 0.66 AS+WC 0.055 0.115 0.07 0.33 0.37 0.42 0.57 0.6 0.6
TABLE IV: Stability of proposed retrieval algorithms when various restrictions are imposed: ”worse–configs (w restr)” – low quality configurations and imposed restrictions; ”best–configs (w restr)” – high quality configurations and imposed restrictions; ”other–configs (w/o restr)” – configurations without restrictions on graph traversal
Additionally, a comparative analysis was conducted, as presented in Table IV, to evaluate the robustness of various graph traversal algorithms. The results indicate that configurations employing 8B models in conjunction with a combined search strategy (A* and WaterCircles) demonstrate high stability, with performance degradation remaining within 4% across varying traversal constraints. In contrast, the BeamSearch algorithm exhibits high sensitivity to these constraints: suboptimal parameterization results in substantial performance loss, with LLM-as-a-Judge varying by as much as 24% between optimal and non-optimal settings. However, in the case of larger-scale models, the combination of BeamSearch and WaterCircles algorithms yields more consistent performance, suggesting improved robustness at higher model capacities.
A critical component of the implemented QA pipeline is the ”NoAnswer” mechanism. This mechanism incorporates a directive in the LLM prompt, instructing the model to output a predefined symbol if provided context lacks sufficient information to generate a valid response. Table V summarizes the frequency of ”NoAnswer” outputs across different model configurations, retrieval algorithms, and graph traversal constraints.
Retrieval
algorithm /
LLM size
WC AS BS WC+BS BS+AS AS+WC
Vertex types
restrictions /
LLM size
all E T O
7B 31% 44% 26% 29% 31% 25% 33% 27% 26% 36%
8B 51% 49% 43% 51% 73% 49% 51% 40% 51% 46%
14B+ 25% 62% 27% 16% 26% 21% 26% 32% 27% 35%
TABLE V: The influence of selected retrieval algorithms and imposed search restrictions on the percentage of generated ”NoAnswer” stubs
As demonstrated in Table V, the analysis reveals distinct patterns in the occurrence of ”NoAnswer” responses across different model configurations. For 7B models, lowest frequency of ”NoAnswer” responses is observed when employing a combined A* and WaterCircles algorithm with restricted traversal of thesis vertices. In contrast, 8B models exhibit minimal ”NoAnswer” instances when utilizing the BeamSearch algorithm alongside a prohibition on episodic vertices traversal. For larger-scale models, the optimal performance, measured by the fewest ”NoAnswer” responses, is achieved through a combined BeamSearch and WaterCircles approach without graph traversal constraints. These findings suggest that the aforementioned algorithms are more effective at extracting relevant information compared to alternative methods under the specified conditions.
It is also important to note the required time to process one user question by QA pipeline with a given LLM and retrieval algorithm: see Table VI.
LLM / Retrieval algorithm QA pipeline Latency (minutes) Mean:
Qwen2.5 7B DeepSeek R1 7B Llama3.1 8B GPT4o mini DeepSeek V3
WC 0.14 0.34 0.46 0.22 0.33 0.30
AS 2.24 4.68 3.51 - 2.53 3.24
BS 5.08 7.86 5.00 8.70 6.32 6.59
TABLE VI: QA pipeline Latency (in minutes), depending on LLM and retrieval algorithm with enabled caching mechanism
Table VI shows that QA pipeline with WaterCircles retriever are significantly faster. This is because that WaterCircles does not need to use vector part of memory graph with vector representations of its elements to perform traversal. Conversely, QA pipeline with BeamSearch algorithm turned out to be slower than QA pipeline with A* algorithm. This is because A* requires embeddings of memory graph elements to construct and traverse only one path on the graph, while BeamSearch requires to construct and monitor $N$ paths/candidates to select optimal traversal. The obtained Latency of the traversal algorithms is also highly dependent on the chosen vector database then configuring vector part of memory. In our experiments Milvus was used. After completing our main experiments, we evaluated read/write performance of five databases (Milvus, OpenSearch, Weaviate, ElasticSearch, and Qdrant) and found the following: (1) Qdrant was the fastest database, and (2) Qdrant in six times faster than Milvus. Therefore, we recommend specifying/using Qdrant when configuring our memory graph to reduce average required time to process single question by our proposed QA pipeline.
In summary, our framework demonstrates that the accuracy can be increased by configuring the memory graph ontology and retrieval methods according to the available LLM and selected QA task. A comparative analysis of existing RAG and GraphRAG methods against our proposed approach, conducted on the TriviaQA and HotpotQA datasets, is provided in Appendix H.
VII Conclusion
This work introduces a flexible and extensible framework for integrating structured memory into language model agents based on a knowledge graph. By extending AriGraph architecture with support of object, episodic, thesis vertices and hyper-adges we enable rich temporal and semantic representations that go beyond traditional RAG pipelines. Our system supports multiple retrieval algorithms, including A*, WaterCircles, BeamSearch and hybrid combinations that can be dynamically adapted to the model’s scale and task requirements. Through extensive evaluation on three benchmarks (DiaASQ, HotpotQA, and TriviaQA), we demonstrate that performance varies systematically with the choice of retrieval strategy and graph traversal constraints. For smaller-scale models (7B–8B), configurations that restrict episodic or object vertices and rely on BeamSearch yield the highest accuracy, while for larger models, hybrid methods combining BeamSearch and WaterCircles offer improved stability and robustness. Importantly, we show that thesis vertices often encode critical information, and excluding them typically degrades performance, especially in 7B models. Compared to existing RAG and GraphRAG methods, our approach demonstrates competitive or superior performance, particularly in handling temporally complex and contradictory information. Ablation studies further reveal that hybrid traversal strategies reduce sensitivity to graph constraints and lower the frequency of invalid responses. Overall, our system provides a principled architecture for long-term, structured memory in LLM agents, enabling personalized, context-aware reasoning at scale. It lays the groundwork for future extensions, involving: temporal filtering, edge-type prioritization and more fine-grained memory control.
VIII Future work
In future work, we propose to enhance the temporal dynamics of our memory graph by introducing a ”memory time” parameter, which will enable fine-grained filtering of triples based on temporal proximity and edge types. This modification will allow the system to selectively prioritize temporally proximate data or emphasize specific relationship categories, thereby improving the precision of personalized responses. Also, one possible future direction is to explore erasure-coded and locally recoverable layouts for sharding graph and vector indices across nodes, inspired by information-theoretic distributed storage, enabling fast repair and continued operation under partial server unavailability [10.1134/S1064226920120116]. In addition, we will investigate private and verifiable retrieval protocols (PIR with result verification) so an agent can query remote memory without revealing the user’s intent and can detect incorrect or malicious responses [10.48550/arXiv.2301.11730]. These mechanisms aim to make long-term personalized memory robust, secure, and auditable at scale.
Appendix A LLM prompts used to build memory graph by Memorize pipeline
Tables VII and VIII presents LLM prompts employed in Memorize pipeline for extracting thesis and object memory triples, respectively, from unstructured natural language text. These prompts facilitate the transformation of textual data into a structured knowledge graph representation.
Type Prompt
System [Uncaptioned image]
User [Uncaptioned image]
Assistant [Uncaptioned image]
TABLE VII: LLM prompts for extracting thesis memories (in the form of triples) from natural language text
Type Prompt
System [Uncaptioned image]
User [Uncaptioned image]
Assistant [Uncaptioned image]
TABLE VIII: LLM prompts for extracting object memories (in the form of triples) from natural language text
Appendix B LLM prompts used to find outdated information in constructed memory graph
Tables IX and X displays LLM prompts employed in Memorize pipeline for identifying stale thesis-related and object memories, respectively, in memory graph.
Type Prompt
System [Uncaptioned image]
User [Uncaptioned image]
Assistant -
TABLE IX: LLM prompts for identifying outdated thesis memories
Type Prompt
System [Uncaptioned image]
User [Uncaptioned image]
Assistant -
TABLE X: LLM prompts for detecting obsolete object memories
Appendix C LLM prompts used in proposed QA pipeline
Table XI presents LLM prompts employed in QA pipeline at the second stage for extracting key entities from the original user question. Table XII displays LLM prompts used in the fourth stage to generate contextually appropriate responses to user question.
Type Prompt
System [Uncaptioned image]
User [Uncaptioned image]
Assistant [Uncaptioned image]
TABLE XI: LLM prompts for extracting key entities from natural language text
Type Prompt
System [Uncaptioned image]
User [Uncaptioned image]
Assistant [Uncaptioned image]
TABLE XII: LLM prompts for conditional generation of answer to user question
Appendix D Preprocessing operations for evaluation datasets
For the original HotpotQA dataset, the distractor/validation subset was selected, comprising 7405 question-answer (QA) pairs and 13781 unique contexts. QA pairs were then filtered to exclude those with associated contexts falling outside a specified length range (measured in characters), retaining only contexts between 64 and 1024 characters in length. This filtering process resulted in 13291 remaining contexts. Finally, first 2000 QA pairs and their corresponding contexts were extracted, yielding a final subset of 3933 unique contexts.
[…]
Thus, evaluation sets for evaluation of proposed/implemented Memorize and QA pipelines were obtained. The characteristics of the obtained subsets of HotpotQA, TriviaQA and DiaASQ datasets can be found in Table XIII.
Dataset QA-pairs Relevant contexts
Amount
Questions length
(in characters)
Answers length
(in characters)
Amount
Length
(in characters)
median mean std median mean std median mean std
DiaASQ 5698 114 109.44 18.66 8 7.57 2.30 3483 556 613.00 324.35
HotpotQA 2000 87 92.98 32.62 13 15.29 11.87 3933 384 413.72 201.05
TriviaQA 500 66 76.37 39.46 9 10.17 5.76 4925 807 765.11 196.32
TABLE XIII: Extended characteristics of datasets, used to evaluate proposed Memorize and QA pipelines
Appendix E Retrieval hyperparameters
  • A*: h_metric_name – ip; max_depth – 10; max_passed_nodes – 150.
  • WaterCircles: strict_filter – True; hyper_num – 15; episodic_num – 15; chain_triplets_num – 25; other_triplets_num – 6; do_text_pruning – False.
  • BeamSearch: max_depth – 5; max_paths – 10; same_path_intersection_by_node – False; diff_paths_intersection_by_node – False; diff_paths_intersection_by_rel – False; mean_alpha – 0.75; final_sorting_mode – mixed.
Appendix F LLM–as–a–Judge instructions
To ensure the reproducibility of the obtained results, LLM inference was conducted using a deterministic generation strategy. The following hyperparameters were applied: num_predict – 2048, seed – 42, temperature – 0.0, and top_k – 1. The Qwen2.5 7B model, sourced from the Ollama repository, was prompted to evaluate whether the outputs of the proposed QA pipeline correctly answered the given questions. The specific LLM prompts used for this assessment are provided in Table XIV.
Type Prompt
System [Uncaptioned image]
User [Uncaptioned image]
Assistant [Uncaptioned image]
TABLE XIV: LLM prompts used in LLM–as–a–Judge framework
Appendix G Characteristics of constructed memory graphs
To evaluate the QA pipeline we constructed 14 memory graphs based on the given dataset and LLM configurations. The structural characteristics of the generated graphs are detailed in Table XV.
Results indicate that certain graphs experienced parsing errors during the extraction of LLM responses, resulting in incomplete storage of contextual information. Across the evaluated datasets, the average parsing error rates were as follows: DiaASQ (7.0%), HotpotQA (6.3%), and TriviaQA (7.3%). For the DiaASQ dataset, DeepSeekR1 7B produced highest number of thesis and object vertices, as well as hyper-edges, while GPT4o-mini extracted largest number of unique object-typed edges. In contrast, for the HotpotQA and TriviaQA datasets, Qwen2.5 7B generated most thesis and object vertices, whereas DeepSeekR1 7B again yielded the highest number of hyper-edges. Based on the obtained results, we can model the structural characteristics of memory graphs generated by each evaluated system: see Figure XVI.
The data reveal significant variation in LLM response parsing accuracy across models. Specifically, Qwen2.5 7B and Llama3.1 8B demonstrate lowest error rates (0.02% each), while DeepSeek V3 exhibits the highest parsing error rate (31.21%). Intermediate performance is observed in DeepSeekR1 7B (0.29%) and GPT4o mini (9.87%). Regarding memory graph composition, DeepSeekR1 7B and Qwen2.5 7B yield the most comprehensive representations, generating the highest number of thesis/object memories and associated edges. Further analysis of vertices creation efficiency per contextual unit shows Qwen2.5 7B achieves superior granularity, producing the largest number of thesis/object memories while maintaining contextual coherence: see Table XVII.
LLM Number of contexts to store in graph Nubmer of vertices Number of edges Average number of matching vertices (by type) episodic thesis object episodic thesis object object neighbours (to episodic vertices) thesis neighbours (to episodic vertices) object neighbours (to thesis vertices) object neighbours (to object vertices) DiaASQ DeepSeek R1 7B 3483 3477 34039 30097 138584 133301 34049 30.06 10.01 3.91 1.99 Qwen2.5 7B 3482 32512 28420 129974 111618 31290 27.98 9.48 3.43 2.04 Llama3.1 8B 3482 29655 20014 100045 77467 28063 20.21 8.52 2.61 2.25 GPT4o mini 3482 31361 28172 125403 96667 39477 27.00 9.00 3.08 1.89 DeepSeek V3 2270 21482 18416 80118 64349 28877 25.83 9,46 2.99 2.11 HotpotQA DeepSeek R1 7B 3933 3921 27137 55254 121359 111174 33714 24.00 7.34 4.09 1.55 Qwen2.5 7B 3933 31795 56078 119387 106037 38460 22.26 8.15 3.33 1.36 Llama3.1 8B 3933 26364 40601 89506 75332 29021 16.04 6.79 2.85 1.47 GPT4o mini 3933 30777 48771 105524 93791 42599 18.99 7.94 3.04 1.36 DeepSeek V3 2713 20164 35291 73242 63921 34621 19.55 7.53 3.17 1.41 TriviaQA DeepSeek R1 7B 4925 4905 48855 96132 213481 201861 52019 33.50 10.56 4.13 1.53 Qwen2.5 7B 4923 52835 109900 220991 188780 62043 34.15 10.87 3.57 1.27 Llama3.1 8B 4922 45241 72285 158202 127389 46757 22.83 9.35 2.81 1.54 DeepSeek V3 3506 37496 68602 143933 122480 61821 30.32 10.75 3.26 1.42
TABLE XV: Characteristics of constructed memory graphs for QA experiments
LLM Number of contexts to store in graph Number of vertices Number of edges Average number of adjacent vertices (by type) episodic thesis object episodic thesis object object neighbours (to episodic vertices) thesis neighbours (to episodic vertices) object neighbours (to thesis vertices) object neighbours (to object vertices) DeepSeek R1 7B 4113 4101 36677 60494 157808 148778 39927 29.18 9.30 4.04 1.69 Qwen2.5 7B 4112 39047 64799 156784 135478 43931 28.13 9.5 3.44 1.55 Llama3.1 8B 4112 33753 44300 115917 93396 34613 19.69 8.22 2.75 1.75 GPT4o mini 3707 31069 38471 115463 95229 41038 22.99 8.47 3.06 1.625 DeepSeek V3 2829 26380 40769 99097 83583 41773 25.23 9.24 3.14 1.64
TABLE XVI: Average/expected characteristics of knowledge graphs in case of using given LLM models for their construction
LLM Vertex Edge
thesis object thesis object
DeepSeek R1 7B 9 15 36 10
Qwen2.5 7B 10 16 33 11
Llama3.1 8B 8 11 23 8
GPT4o mini 8 10 26 11
DeepSeek V3 9 14 30 15
TABLE XVII: Number of unique vertices/edges that were added to the graph when processing (using a given LLM model) and storing one episodic memory (episodic verrtex)
In addition to characteristics of constructed graphs, we collected information about time and speed of the Memorize pipeline, which responsible for parsing incoming unstructured texts in natural language and storing them in memory model: see Table XVIII.
LLM / Dataset Memory construction time (hours) and speed (contexts per minute) Mean:
Qwen2.5 7B DeepSeek R1 7B Llama3.1 8B GPT4o mini DeepSeek V3
DiaASQ 23.5 / 2.47 57.5 / 1.00 32 / 1.81 32 / 1.81 68 / 0.85 42.6 / 1.58
HotpotQA 61.5 / 1.06 47.5 / 1.38 27 / 2.42 34 / 1.92 72 / 0.91 48.4 / 1.53
TriviaQA 90 / 0.91 - 90 / 0.91 - 80 / 1.02 86.6 / 0.94
Mean: 58.3 / 1.48 52.5 / 1.19 49.7 / 1.71 33 / 1.86 73.3 / 0.92
TABLE XVIII: Time and speed of memory graphs construction based on a given dataset and LLM, which were used in proposed QA pipeline evaluation
Table XVIII shows that highest speed of parsing and saving information into memory is observed in configurations of Memorize pipeline there GPT4o mini is set as LLM (1.86); in second place is LLama3.1 8B, that was deployed in Ollama Docker container on local machine.
It is also important to note the amount of disk space that constructed memory is required. When using Milvus as a database for storing vector representations of text, on selected datasets, one memory graph in our implementation occupies, approximately, 80-90 GB. If Qdrant is used as a vector storage, the same graphs will occupy, approximately, 4-6 GB each.
Appendix H Comparison with existing RAG and GraphRAG methods
Based on experimental results, a comparative table, summarizing best-performing QA configurations by Exact Match metric, was compiled (see Table XIX).
LLM /
Dataset
Qwen2.5 7B DeepSeek R1 7B Llama3.1 8B GPT4o mini DeepSeek V3
DiaASQ 0.22 / BS + WC / T 0.11 / AS / E 0.18 / AS / E 0.47 / BS + WC / E 0.46 / BS + WC / O
HotpotQA 0.18 / AS / all 0.14 / AS / E 0.37 / BS / O 0.59 / BS + WC / all 0.6 / BS + WC / T
TriviaQA 0.2 / BS / E 0.18 / AS / E 0.47 / BS / E - 0.62 / BS + WC / all
Mean: 0.2 0.14 0.34 0.53 0.56
TABLE XIX: Best QA configurations ranked by Exact Match metric across all experiments. The corresponding cells contains Exact Match score, used retrieval algorithm and the type of restriction applied to the memory graph during traversal. Shortcuts for retrieval algorithms: BA – BeamSearch; AS – A*; BS + AS – hybrid of BeamSearch and A*; BS + WC – hybrid of BeamSearch and WaterCircles. Shortcuts for graph restrictions: all – no restrictions applied; E – episodic vertices excluded from traversal; T – thesis vertices excluded; O – object vertices excluded.
As shown in Table XIX, Qwen2.5 achieve best performance (0.2) among the 7B models. Among all evaluated configurations, highest overall effectiveness (0.56) was reached by setups incorporating DeepSeek V3. Notably, top-performing 7B configurations predominantly relied on A*, especially under constraint that traversal restricted through episodic vertices. In contrast, best DeepSeek V3 configurations frequently adopted a hybrid strategy combining BeamSearch and WaterCircles. Across high-performing configurations more broadly, BeamSearch consistently appeared as a key component within the retrieval pipeline.
To evaluate the performance of optimal configurations derived from our framework against existing Retrieval-Augmented Generation (RAG) and GraphRAG approaches, we conducted a systematic literature review. The search was performed across five academic search engines: SciSpace, Scite, PaperDigest, Consensus, and Elicit. Using each engine, we selected the first ten publications in the search results for the following queries: in case of RAG methods – ”Retrieval-augmented generation methods based on pretrained language models” and ”RAG methods in NLP”; in case of GraphRAG methods – ”RAG on Knowledge Graphs”, ”Enhancing RAG–approach with Knowledge Graphs”, ”Graph RAG” and ”RAG with integration of Large Language Models (LLMs) and Knowledge Graphs (KGs)”. The search was restricted to publications up to 2018. Subsequently, we applied a three-stage filtering process: (1) duplicate entries were removed, and papers introducing novel evaluation datasets were excluded; (2) works focusing on domain-specific applications without broader methodological contributions were discarded; (3) only studies providing comprehensive methodological descriptions and employing standardized benchmarks were retained. This process yielded two final sets of nine articles each, covering RAG and GraphRAG techniques, respectively. A comparative analysis of these methods against our framework’s optimal graph construction and retrieval configuration is presented in Table XX, with performance measured using ExactMatch metric.
Dataset RAG method Our method
REALM [10.5555/3524938.3525306] DPR [karpukhin-etal-2020-dense] RAG [10.5555/3495724.3496517] ColBERT–QA [khattab-etal-2021-relevance] FiD [izacard-grave-2021-leveraging] EMDR2 [sachan2021endtoend] RETRO [Borgeaud2021ImprovingLM] Atlas [10.5555/3648699.3648950] RePLUG [shi-etal-2024-replug]
TriviaQA 53.9 56.8 55.8 70.1 67.6 71.4 62.1 79.8 77.3 62.0
GraphRAG method
ToG [sun2023thinkongraph] RoG [luo2024rog] PMKGE [liu2025enhancinglargelanguagemodels] GRAG [hu2024grag] GNN–RAG [mavromatis-karypis-2025-gnn] ToG2.0 [Ma2024ThinkonGraph2D] DoG [ma2025debate] GCR [luo2024graph] PDA [Sun2024PyramidDrivenAP]
HotpotQA 41,0 43.0 42.6 36.1 43.0 40.9 45.3 45.9 36.5 60.0
TABLE $X$: Comparison of existing RAG and GraphRAG methods with our proposed method on ExactMatch metric. On TriviaQA dataset QA configuration with DeepSeek V3 and combination of BeamSearch and WaterCircles algorithms (without restrictions on graph traversal) were used. On HotpotQA dataset QA configuration with GPT4o–mini and the same algorithm and graph restrictions were used.
As shown in Table XX, our proposed method achieves a 14.1% improvement over existing GraphRAG approaches in a specific configuration. However, it underperforms compared to standard RAG methods by 17.8%. This discrepancy can be attributed to the fact that the RAG baselines evaluated in this study employed Reader and Retriever models that were specifically fine-tuned on the same dataset used for evaluation. As established in prior works, such in-domain fine-tuning typically yields optimal performance, and evaluation on out-of-domain datasets would be expected to result in significant degradation of measured metrics.
We also reproduce and evaluate the quality of HippoRAG method on DiaASQ and HotpotQA datasets: DiaASQ – 0.53 (LLM-as-a-Judge); HotpotQA – 60.2 (Exact Match). It can be seen that our method shows comparable results or outperforms HippoRAG.
[…]